OpenStudy (anonymous):

how to factor 5x^2 +4x -1

OpenStudy (anonymous):

Step by step would be nice :) also some tips on factoring

OpenStudy (anonymous):

(5x )(x ) 5x times 1x = 5x^2 (5x 1)(x 1) 1 times 1 = 1 Now what are the signs? if it was (5x+1)(x-1), it would be -5x+x, which does not equal 4x However, if it was (5x-1)(x+1), it would be 5x+(-x) which equals 4x. Therefore, (5x-1)(x+1) is your answer.

OpenStudy (anonymous):

You have a Quadratic equation so the best thing to do is to use the Quadratic Formula. The Quadratic formula is \[-b \pm \sqrt{b^2 - 4ac}/2a\] So in the problem 5x^2 +4x -1 a = 5 b = 4 c = -1 plug in the numbers into the formula and you get \[-4 \pm \sqrt{16 - 4(5)(-1)}/10\] (note the whole thing is divided by 10 not just the radical). Simplify the expression inside the radical and you get \[-4 \pm \sqrt{36}/10\] Evaluate the square root and you get \[-4 \pm 6 /10\] So your two solutions are \[-4 -6 / 10\] and \[-4 + 6 /10\] Simplify and you get\[-10/10\] and \[2/10\] Which simplify down to -1 and 1/5

OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

5x^(2)+4x-1 For a polynomial of the form ax^(2)+bx+c, find two factors of a*c (-5) that add up to b (4). a=5, b=4, c=-1 For a polynomial of the form ax^(2)+bx+c, find two factors of a*c (-5) that add up to b (4).In this problem 1*-(1)/(5)=-(1)/(5) (which is (c)/(a)) and 1-(1)/(5)=(4)/(5) (which is (b)/(a)) , so insert 1 as the right hand term of one factor and -(1)/(5) as the right-hand term of the other factor. (x+1)(x-(1)/(5)) Remove the fraction by multiplying the first term of the factor by the denominator of the second term. (x+1)(5x-1)